# Incremental Generation is Necessity and Sufficient for Universality in Flow-Based Modelling

### v1

@@ B.1.2 Differentiable Case
$\begin{split}\big\|\varphi(x)-\bigcirc_{g=1}^{G}\,\Psi^{(g)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}&=\big\|\bigcirc_{g=1}^{G}\varphi^{(g)}(x)-\bigcirc_{g=1}^{G}\,\Psi^{(g)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\\ &\leq\big\|\bigcirc_{g=1}^{G}\,\varphi^{(g)}(x)-\varphi^{(G)}\circ\bigcirc_{g=1}^{G-1}\Psi^{(g)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\\ &\quad\quad+\big\|\varphi^{(G)}\circ\bigcirc_{g=1}^{G-1}\Psi^{(g)}(x)-\bigcirc_{g=1}^{G}\,\Psi^{(g)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\\ &\leq\big\|\bigcirc_{g=1}^{G-1}\varphi^{(g)}(x)-\bigcirc_{g=1}^{G-1}\,\Psi^{(g)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}e^{L^{V^{{(G)}}}}\\ &\quad\quad+2\|\omega^{(G)}(N_{1},\dots,N_{d},L)\|_{l^{\infty}}e^{L^{V^{{(G)}}}}\end{split}$
@@
$\big\|\varphi(x)-\bigcirc_{g=1}^{G}\,\Psi^{(g)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\sum_{g=1}^{G}\left(2\|\omega^{(g)}(N_{1},\dots,N_{d},L)\|_{l^{\infty}}\prod_{j=g}^{G}e^{L^{V^{(j)}}}\right)$
@@ B.1.3 Smooth Case
Lemma 23 (Large Granularity Triviality for Smooth Regularity)
There exists a $K_{d}\in\mathbb{N}_{+}$ s.t. for any integer $G>K_{d}$, $\operatorname{\mathcal{H}}_{d}^{G,\infty}([0,1]^{d})$ is empty.
Proof : Note $\operatorname{\mathcal{H}}_{d}^{G,\infty}\left([0,1]^{d}\right)=\operatorname{\mathcal{H}}_{d}^{G,\infty}\left((0,1)^{d}\right)$ because these set of functions are zero on the boundary of $0,1]^{d}$ by continuity. Define the autonomous norm
$\|f\|_{\mathrm{frag}}:=\min\left\{m\in\mathbb{N}\mid f=h_{1}\cdots h_{m}\text{ where }h_{i}=\mathrm{Flow}(V_{i})\text{ for some }V_{i}\in\mathcal{X}\right\}$
@@
For every diffeomorphism $\varphi\in\operatorname{Diff}_{0}([0,1]^{d})$, there exist some $G\in\mathbb{N}$ and flows $\varphi^{(1)},\dots,\varphi^{(G)}\in\operatorname{Flow}{}$ such that
$\varphi=\bigcirc_{g=1}^{G}\,\varphi^{(i)}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\varphi^{(G)}\circ\dots\circ\varphi^{(1)}.$ (33)
@@
Proof of Theorem 9: By Lemma 24 there exists a $G$ such that $\varphi\in\operatorname{\mathcal{H}}_{d}^{G,\infty}([0,1]^{d})$; as $\operatorname{\mathcal{H}}_{d}^{G,\infty}([0,1]^{d})$ is non-empty, by Lemma 23 $G\leq K_{d}$. This means one can write any diffeomorphism as a composition of at most $K_{d}$ flows.
@@
Proof of Proposition 10: As $\varphi\in\operatorname{Diff}_{0}([0,1]^{d})$ it is also $s$ differentiable so by Theorem 9 there exists $G\leq K_{d}$ such that $\varphi\in\operatorname{\mathcal{H}}_{d}^{G,s}([0,1]^{d})$. By Proposition 8 there exists $\Psi$ such that:
$\begin{split}\big\|\varphi-\Psi\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}&\leq\sum_{g=1}^{G}\left(2\|\omega^{(g)}(N,L)\|_{l^{\infty}}\prod_{j=g}^{G}e^{L^{V^{(j)}}}\right)\\ &\leq 170(s+1)^{d}8^{s}\max_{g,j}\{\|V^{(g)}_{j}\|_{C^{s}([0,1]^{d})}\}(NL)^{-2s/d}e^{\max_{g}\{L^{V^{(g)}}\}}K_{d}\left(\prod_{j=2}^{G}L^{V^{(j)}}\right)\\ &\in O\big(N^{-2s/d}\big)\end{split}$ (34)
as $(\omega^{(g)})_{j}=85(s+1)^{d}8^{s}\|V^{(g)}_{j}\|_{C^{s}([0,1]^{d})}(NL)^{-2s/d}$
@@
We may without loss of generality consider $\mathrm{sup}(\varphi\circ\varphi_{M}^{-1})=B(0,\frac{1}{8})$ thus $\varphi\circ\varphi_{M}^{-1}$ fixes all points on a neighborhood of the boundary of $B(0,1/4)$. Consequently, we may apply the Munkres-Connel-Bing Theorem, as formulated in (muller2014uniform, Lemma 2) to deduce that: for every $\varepsilon>0$ there exists a diffeomorphism $\varphi_{\varepsilon}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ supported on $B(0,\frac{1}{4})$ satisfying the uniform approximation guarantees
@@
Now, since $\varphi_{\varepsilon}$ is supported on $B(0,\frac{1}{4})\subset[0,1]^{d}$ and since $\varphi_{\varepsilon}$ is a diffeomorphism then Theorem 9 implies that there exists some $\tilde{K}_{d}$ (not depending on $\varphi_{\varepsilon}$ nor on $\varepsilon$) and some $1\leq G\leq\tilde{K}_{d}$ such that $\varphi_{\varepsilon}\in\operatorname{\mathcal{H}}_{d}^{G,1}([0,1]^{d})$. Applying Theorem 7 implies that there exists ReLU neural ODEs $\{\Psi^{(g)}=\mathrm{Flow}(\Phi^{(g)})\}_{g=1}^{G}\subset\mathcal{X}^{0}$ such that the diffeomorphism $\Psi_{\varepsilon}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{g=1}^{G}\,\Psi^{(g)}$ is Lipschitz and compactly supported and satisfies the approximation guarantee:
@@
and setting $\Psi=\Psi_{\varepsilon}\circ\Psi_{M}$ and $K_{d}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\tilde{K}_{d}+1$ yields the conclusion.
Appendix C Additional Geometric and Topological Background
@@
    Nowhere dense and meagerness: Let $T=(S,\tau)$ be a topological space and $\subseteq S. AA$ is nowhere dense in $T$ if and only if
@@
    AA is meager in $T$ if and only if it can be written as a countable union
@@
Appendix D Proofs of Corollaries
$\|f-\Psi\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}<\epsilon.$
also width and depth are given in table 1. Also note that by lifting, projecting and concatination the function remains Lipschitz. This completes our proof.
Proof of Corollary 14:
Fix $\varepsilon>0$, applying Corollary 11 we find that there exists a Lipschitz Latent Neural ODE $\Psi:\mathbb{R}^{d+D}\to\mathbb{R}^{d+D}$ induced by ReLU MLP of depth $\mathcal{O}(\log_{2}(d))$, width $\mathcal{O}(dn^{d+1})$, with $\mathcal{O}(dn^{d+1})$ non-zero parameters, such that
So
It remains to control term (II). In particular,
But
Consequently, combining (40), (44) ,and (47) yields the conclusion.

### v2

@@ B.1.2 Differentiable Case
$\begin{split}\big\|\varphi(x)-\bigcirc_{t=1}^{T}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}&=\big\|\bigcirc_{t=1}^{T}\varphi^{(t)}(x)-\bigcirc_{t=1}^{T}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\\ &\leq\big\|\bigcirc_{t=1}^{T}\,\varphi^{(t)}(x)-\varphi^{(T)}\circ\bigcirc_{t=1}^{T-1}\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\\ &\quad\quad+\big\|\varphi^{(T)}\circ\bigcirc_{t=1}^{T-1}\Psi^{(t)}(x)-\bigcirc_{t=1}^{T}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\\ &\leq\big\|\bigcirc_{t=1}^{T-1}\varphi^{(t)}(x)-\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}e^{L^{V^{{(T)}}}}\\ &\quad\quad+2\|\omega^{(T)}(N_{1},\dots,N_{d},L)\|_{l^{\infty}}e^{L^{V^{{(T)}}}}\end{split}$
@@
$\big\|\varphi(x)-\bigcirc_{t=1}^{T}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\sum_{t=1}^{T}\left(2\|\omega^{(t)}(N_{1},\dots,N_{d},L)\|_{l^{\infty}}\prod_{j=t}^{T}e^{L^{V^{(j)}}}\right)$
@@ B.1.3 Smooth Case
Lemma 24 (Large incrementality Triviality for Smooth Regularity)
For $\delta>0$, there exists $T_{d}\in\mathbb{N}_{+}$ s.t. for any integer $T>T_{d}$, $\operatorname{\mathcal{H}}_{d}^{T,\infty}((-\delta,1+\delta)^{d})$ is empty.
Proof : Define $S:=\{\,\operatorname{Flow}{}(V)\mid V\text{ smooth, compactly supported in }(0,1)^{d}\,\}$ and the norm
$\|f\|_{\mathrm{frag}}=\begin{cases}0,&f=\mathrm{id},\\[6.0pt] \displaystyle\min\{\,m\in\mathbb{N}\mid f=h_{1}\cdots h_{m},\;h_{i}\in S\,\},&f\neq\mathrm{id}.\end{cases}$
@@
For every diffeomorphism $\varphi\in\operatorname{Diff}_{0}([0,1]^{d})$, there exist some $T\in\mathbb{N}$ and flows $\varphi^{(1)},\dots,\varphi^{(T)}\in\operatorname{Flow}{}$ such that
$\varphi=\bigcirc_{t=1}^{T}\,\varphi^{(i)}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\varphi^{(T)}\circ\dots\circ\varphi^{(1)}.$ (33)
@@
Proof of Theorem 10: By Lemma 25 there exists a $T$ such that $\varphi\in\operatorname{\mathcal{H}}_{d}^{T,\infty}([0,1]^{d})$; as $\operatorname{\mathcal{H}}_{d}^{T,\infty}([0,1]^{d})$ is non-empty and for $0<\delta$ we have $\operatorname{\mathcal{H}}_{d}^{T,\infty}([0,1]^{d})\subset\operatorname{\mathcal{H}}_{d}^{T,\infty}((-\delta,1+\delta)^{d})$, by applying Lemma 24 we conclude $T\leq T_{d}$. This means one can write any diffeomorphism as a composition of at most $T_{d}$ flows.
@@
Proof of Proposition 11: As $\varphi\in\operatorname{Diff}_{0}([0,1]^{d})$ it is also $s$ differentiable so by Theorem 10 there exists $T\leq T_{d}$ such that $\varphi\in\operatorname{\mathcal{H}}_{d}^{T,s}([0,1]^{d})$. By Proposition 9 there exists $\Psi$ such that:
$\begin{split}\big\|\varphi-\Psi\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}&\leq\sum_{t=1}^{T}\left(2\|\omega^{(t)}(N,L)\|_{l^{\infty}}\prod_{j=t}^{T}e^{L^{V^{(j)}}}\right)\\ &\leq 170(s+1)^{d}8^{s}\max_{t,j}\{\|V^{(t)}_{j}\|_{C^{s}([0,1]^{d})}\}(NL)^{-2s/d}e^{\max_{t}\{L^{V^{(t)}}\}}T_{d}\left(\prod_{j=2}^{T}L^{V^{(j)}}\right)\\ &\in O\big(N^{-2s/d}\big)\end{split}$ (34)
as $(\omega^{(t)})_{j}=85(s+1)^{d}8^{s}\|V^{(t)}_{j}\|_{C^{s}([0,1]^{d})}(NL)^{-2s/d}$
B.2 Qualitative Universal Approximation Guarantee
@@
We may without loss of generality consider $\mathrm{supp}^{\dagger}(\varphi\circ\varphi_{M}^{-1})=B(0,\frac{1}{8})$ thus $\varphi\circ\varphi_{M}^{-1}$ fixes all points on a neighbourhood of the boundary of $B(0,1/4)$. Consequently, we may apply the Munkres-Connel-Bing Theorem, as formulated in (muller2014uniform, Lemma 2) to deduce that: for every $\varepsilon>0$ there exists a diffeomorphism $\varphi_{\varepsilon}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ supported on $B(0,\frac{1}{4})$ satisfying the uniform approximation guarantees
@@
Now, since $\varphi_{\varepsilon}$ is supported on $B(0,\frac{1}{4})\subset[0,1]^{d}$ and since $\varphi_{\varepsilon}$ is a diffeomorphism then Theorem 10 implies that there exists some $\tilde{T}_{d}$ (not depending on $\varphi_{\varepsilon}$ nor on $\varepsilon$) and some $1\leq T\leq\tilde{T}_{d}$ such that $\varphi_{\varepsilon}\in\operatorname{\mathcal{H}}_{d}^{T,1}([0,1]^{d})$. Applying Theorem 8 implies that there exists ReLU neural ODEs $\{\Psi^{(t)}=\mathrm{Flow}(\Phi^{(t)})\}_{t=1}^{T}\subset\mathcal{X}^{0}$ such that the homeomorphism $\Psi_{\varepsilon}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{t=1}^{T}\,\Psi^{(t)}$ is Lipschitz and compactly supported and satisfies the approximation guarantee:
@@
and setting $\Psi=\Psi_{\varepsilon}\circ\Psi_{M}$ and $T_{d}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\tilde{T}_{d}+1$ yields the conclusion. [^9] Note that because $\varphi_{M}\in\operatorname{\mathcal{H}}_{d}([-M,M]^{d})$ we have $\mathrm{supp}^{\dagger}(\Psi)=[-M,M]^{d}$
Appendix C Proofs of Corollaries
$\|f-\Psi\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}<\epsilon.$
also width and depth are given in table 1. Also note that by lifting, projecting and concatenation the function remains Lipschitz. This completes our proof.
Proof of Corollary 15:
Fix $\varepsilon>0$, applying Corollary 12 we find that there exists a Lipschitz Latent Neural ODE $\Psi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ induced by ReLU MLP of depth $\mathcal{O}(\log_{2}(d))$, width $\mathcal{O}(dn^{d+1})$, with $\mathcal{O}(dn^{d+1})$ non-zero parameters, such that
So
It remains to control term (II). In particular,
But
Consequently, combining (40), (44) ,and (47) yields the conclusion.
Appendix D Additional Geometric and Topological Background
@@
    Nowhere dense and meagreness: Let $T=(S,\tau)$ be a topological space and $\subseteq S. AA$ is nowhere dense in $T$ if and only if
@@
    AA is meagre in $T$ if and only if it can be written as a countable union